# A Numerical Study on Gauge Symmetry of Electroweak Amplitudes

### v2

@@ V Anomalous Couplings and SMEFT
So far the anomalous couplings we considered are only instruments to demonstrate the gauge symmetry of related amplitudes. They do not have intrinsic, gauge-invariant, physical meanings. However, there is a scenario that anomalous couplings can be physical. That is when those couplings are part of a set of anomalous SM couplings from new physics that respects gauge symmetry. An example that we will explore in this section is SMEFT, which is an effective field theory respecting the SM gauge symmetry. In particular, we will discuss two dim-6 SMEFT operators Grzadkowski:2010es ; Dedes:2017zog ; Dedes:2019uzs ; Barger:2023wbg : $\mathcal{O}_{6}=C_{6}(\Phi^{\dagger}\Phi)^{3}$ and $\mathcal{O}_{t\Phi}=C_{t\Phi}(\Phi^{\dagger}\Phi)(\bar{Q}_{\mathrm{L}}t_{\mathrm{R}}\tilde{\Phi})+\mathrm{H.c.}$, where $\Phi$ is the SM Higgs doublet and $\tilde{\Phi}\equiv i\sigma^{2}\Phi^{*}$. We will study how the gauge symmetry is restored in the presence of these operators when the 3-point Higgs self-coupling $\lambda_{hhh}$ and the top Yukawa coupling $y_{tth}$ are modified in the $WW\rightarrow hh$ and $WW\rightarrow t\bar{t}$ processes, respectively.
@@
Adding an SMEFT operator $\mathcal{O}_{6}=C_{6}(\Phi^{\dagger}\Phi)^{3}$ into the SM Lagrangian, the $hhh$ and $hh\varphi^{+}\varphi^{-}$ couplings are modified to
@@
Since the two couplings are modified by the same operator, their deviations from the SM values are related. This relation is further protected by gauge symmetry. Conversely, if we modify one of the couplings alone, say $\lambda_{hhh}=\lambda_{hhh}^{\mathrm{SM}}+\delta_{hhh}$, the gauge symmetry will be broken. In order to restore gauge symmetry, the other coupling $\lambda_{hh\varphi^{+}\varphi^{-}}$ has to be modified by adding
@@
Any deviation from the above relation will break gauge symmetry for a process involving both $hhh$ and $hhVV$ couplings, which can be tested with the MWI.
In order to test the relation (20), we introduce a deviation parameter $\delta$ by
@@
Thus, $\delta=0$ corresponds to the relation (20) that respects gauge symmetry. We take $\delta_{hhh}$ to be 0.20.20.2, 0.40.40.4, 0.80.80.8, and 111. For every $\delta_{hhh}$ value, we compute $k_{M}\mathcal{M}^{M}$ as a function of $\delta$ for the $WW\rightarrow hh$ process with one $W$ polarization vector replaced by the 5-component momentum. The results are summarized in Fig. 6. As can be seen in the figure, $k_{M}\mathcal{M}^{M}=0$ only appears when $\delta=0$, i.e., when the relation (20) holds. This confirms our argument that, with an anomalous coupling of $\lambda_{hhh}$, gauge symmetry can be restored by adding an proper anomalous coupling of $\lambda_{hh\varphi\varphi}$.
Figure 6: $k_{M}\mathcal{M}^{M}$ for $WW\rightarrow hh$ as functions of $\delta$ with different values of $\delta_{hhh}$, serving as a test for gauge symmetry in the presence of the SMEFT operator $\mathcal{O}_{6}$.
@@
We then focus on gauge symmetry of the $WW\rightarrow t\bar{t}$ process with the incorporation of the SMEFT operator $\mathcal{O}_{t\Phi}=C_{t\Phi}(\Phi^{\dagger}\Phi)(\bar{Q}_{\mathrm{L}}t_{\mathrm{R}}\tilde{\Phi})+\mathrm{H.c.}$. The couplings of $tth$ and $tt\varphi\varphi$ are defined as
@@
After the addition of $\mathcal{O}_{t\varPhi}$, the couplings are modified into
@@
Following the above equations, we can obtain the relations between the modifications to these couplings are
@@
If expressing in left-handed and right-handed couplings, the Lagrangian terms for the $\bar{t}th$ and $tt\varphi\varphi$ couplings become
@@
It is easy to see that they are related to each other by
@@
Our method of testing the gauge symmetry is similar to $WW\rightarrow hh$ with the $\mathcal{O}_{6}$ operator. Two deviation parameters $\delta_{\mathrm{L}}$ and $\delta_{\mathrm{R}}$ are introduced by
@@
Therefore, $\delta_{\mathrm{L}}=\delta_{\mathrm{R}}=0$ corresponds to the modifications contributed by the $\mathcal{O}_{t\Phi}$ operator that preserve gauge symmetry. In Figs. LABEL:fig:smeft_ww2tt(a) and LABEL:fig:smeft_ww2tt(b), we show $k_{M}\mathcal{M}^{M}$ for $WW\rightarrow t\bar{t}$ as functions of $\delta_{\mathrm{L}}$ and $\delta_{\mathrm{R}}$ assuming various values of $\operatorname{Re}(\delta_{tth})$ and $\operatorname{Im}(\delta_{tth})$, where $\delta_{tth}$ is defined as $\delta_{tth}\equiv C_{t\Phi}v^{2}/\sqrt{2}$. As expected, gauge symmetry is restored when the condition (30) is satisfied. This demonstrates that an anomalous Yukawa coupling can be counterbalanced by a corresponding $ff\varphi\varphi$ contact vertex, ensuring that the $WW\rightarrow t\bar{t}$ amplitude remains gauge invariant.
@@ VI Summary and Conclusions
In this paper, we demonstrated that the GE representation of EW interactions, which we had an brief introduction first, has gauge symmetry imprinted in the structure of amplitudes, manifesting in the MWI (1). We numerically studied this important property of EW interactions in the GE representation in many different aspects, including directly testing the MWI, modifying couplings within vertices, and modifying overall couplings of vertices.
@@
By directly testing gauge symmetry on 4-point amplitudes such as $WW\rightarrow t\bar{t}$ and $W^{+}W^{-}\rightarrow W^{+}W^{-}$ with different helicity combinations, we found that the full amplitudes always satisfy the MWI after summing over all tree-level diagrams. This result typically involves large cancellations between individual diagrams, similar to unitarity cancellation in the gauge representation.
@@
For testing gauge symmetry on 3-point vertices with anomalous couplings, we studied the $VVh$, $ff^{\prime}V$, and $VVV$ vertices. We found that there are precise relations governing the couplings of Goldstone and gauge components for these vertices. We then numerically test if the MWI still holds when some of the couplings are modified so that those relations are violated. What we found is that any deviation from those relations by anomalous couplings would violate the MWI.
@@
For testing gauge symmetry on 4-point amplitudes with anomalous couplings, we studied $W^{+}W^{+}\rightarrow W^{+}W^{+}$, $WW\rightarrow hh$, and $WW\rightarrow t\bar{t}$. We focused on how modifying the overall couplings of individual vertices affects gauge symmetry. Our results are similar to 3-point vertices: gauge symmetry is manifested as precise relations between couplings, and modifying couplings to violate those relations also results in violating the MWI.
@@
After studying gauge symmetry of the SM, we then proceed to study effective operators in the SMEFT, specifically the $\mathcal{O}_{6}$ operator that modifies the Higgs self-couplings and the $\mathcal{O}_{t\Phi}$ operator that modifies the top Yukawa coupling. We found that the operators modify both the couplings mentioned and the related Goldstone couplings, giving relations between the couplings involving $C_{6}$ and $C_{t\Phi}$, respectively. Numerically testing the MWI indicates that gauge symmetry is preserved as long as all related couplings are modified in accordance with those relations. On the other hand, if those relations are violated, gauge symmetry would be broken.
@@
We believe that this paper contributes to a deeper understanding of EW interactions and massive gauge theory in general. In addition, it also gives a convenient way of checking self-consistency of EW interactions in the GE representation.

### v3

@@ V Anomalous Couplings and SMEFT
Thus far, the anomalous couplings we considered are only instruments to demonstrate the gauge symmetry of related amplitudes. They do not have intrinsic, gauge-invariant, physical meanings. However, there is a scenario in which anomalous couplings can be physical: when those couplings are part of a set of anomalous SM couplings from new physics that respects gauge symmetry. An example that we explore in this section is the SMEFT, which is an effective field theory respecting the SM gauge symmetry. In particular, we discuss two dim-6 SMEFT operators Grzadkowski:2010es ; Dedes:2017zog ; Dedes:2019uzs ; Barger:2023wbg : $\mathcal{O}_{6}=C_{6}(\Phi^{\dagger}\Phi)^{3}$ and $\mathcal{O}_{t\Phi}=C_{t\Phi}(\Phi^{\dagger}\Phi)(\bar{Q}_{\mathrm{L}}t_{\mathrm{R}}\tilde{\Phi})+\mathrm{H.c.}$, where $\Phi$ is the SM Higgs doublet, and $\tilde{\Phi}\equiv i\sigma^{2}\Phi^{*}$. We study how the gauge symmetry is restored in the presence of these operators when the three-point Higgs self-coupling $\lambda_{hhh}$ and the top Yukawa coupling $y_{tth}$ are modified in the $WW\rightarrow hh$ and $WW\rightarrow t\bar{t}$ processes, respectively.
@@
Adding an SMEFT operator $\mathcal{O}_{6}=C_{6}(\Phi^{\dagger}\Phi)^{3}$ into the SM Lagrangian, the hhhhhh and $hh\varphi^{+}\varphi^{-}$ couplings are modified to
@@
As the two couplings are modified by the same operator, their deviations from the SM values are related. This relation is further protected by gauge symmetry. Conversely, if we modify one of the couplings alone, say $\lambda_{hhh}=\lambda_{hhh}^{\mathrm{SM}}+\delta_{hhh}$, the gauge symmetry will be broken. To restore gauge symmetry, the other coupling $\lambda_{hh\varphi^{+}\varphi^{-}}$ has to be modified by adding
@@
Any deviation from the above relation will break gauge symmetry for a process involving both hhhhhh and $hhVV$ couplings, which can be tested with the MWI.
To test the relation (20), we introduce a deviation parameter $\delta$ as follows:
@@
Thus, $\delta=0$ corresponds to the relation (20), which respects gauge symmetry. We consider $\delta_{hhh}$ to be $0.2$, $0.4$, $0.8$, and $1$. For every $\delta_{hhh}$ value, we compute $k_{M}\mathcal{M}^{M}$ as a function of $\delta$ for the $WW\rightarrow hh$ process with one $W$ polarization vector replaced by the five-component momentum. The results are summarized in Fig. 6. As shown in the figure, $k_{M}\mathcal{M}^{M}=0$ only appears when $\delta=0$, i.e., when the relation (20) holds. This confirms our argument that, with an anomalous coupling of $\lambda_{hhh}$, gauge symmetry can be restored by adding a proper anomalous coupling of $\lambda_{hh\varphi\varphi}$.
Figure 6: $k_{M}\mathcal{M}^{M}$ for $WW\rightarrow hh$ as a function of $\delta$ with different values of $\delta_{hhh}$, serving as a test for gauge symmetry in the presence of the SMEFT operator $\mathcal{O}_{6}$.
@@
We then focus on the gauge symmetry of the $WW\rightarrow t\bar{t}$ process with the incorporation of the SMEFT operator $\mathcal{O}_{t\Phi}=C_{t\Phi}(\Phi^{\dagger}\Phi)(\bar{Q}_{\mathrm{L}}t_{\mathrm{R}}\tilde{\Phi})+\mathrm{H.c.}$. The couplings of tthtth and $tt\varphi\varphi$ are defined as
@@
After the addition of $\mathcal{O}_{t\varPhi}$, the couplings are modified to
@@
Based on the above equations, we can obtain the relations between the modifications to these couplings as
@@
If expressed in left-handed and right-handed couplings, the Lagrangian terms for the $\bar{t}th$ and $tt\varphi\varphi$ couplings become
@@
We can observe that they are related to each other by
@@
Our method of testing the gauge symmetry is similar to that for $WW\rightarrow hh$ with the $\mathcal{O}_{6}$ operator. Two deviation parameters $\delta_{\mathrm{L}}$ and $\delta_{\mathrm{R}}$ are introduced as follows:
@@
Therefore, $\delta_{\mathrm{L}}=\delta_{\mathrm{R}}=0$ corresponds to the modifications contributed by the $\mathcal{O}_{t\Phi}$ operator that preserve gauge symmetry. In Figs. 8(a)(a) and 8(a)(b), we show $k_{M}\mathcal{M}^{M}$ for $WW\rightarrow t\bar{t}$ as functions of $\delta_{\mathrm{L}}$ and $\delta_{\mathrm{R}}$ assuming various values of $\operatorname{Re}(\delta_{tth})$ and $\operatorname{Im}(\delta_{tth})$, where $\delta_{tth}$ is defined as $\delta_{tth}\equiv C_{t\Phi}v^{2}/\sqrt{2}$. As expected, gauge symmetry is restored when the condition in Eq. (30) is satisfied. This demonstrates that an anomalous Yukawa coupling can be counterbalanced by a corresponding $ff\varphi\varphi$ contact vertex, ensuring that the $WW\rightarrow t\bar{t}$ amplitude remains gauge invariant.
(a) $k_{M}\mathcal{M}^{M}$ for $WW\rightarrow t\bar{t}$ as a function of $\delta_{\mathrm{L}}$ (a) and $\delta_{\mathrm{R}}$ (b) with different values of $\operatorname{Re}(\delta_{tth})$ and $\operatorname{Im}(\delta_{tth})$, serving as a test for gauge symmetry in the presence of the SMEFT operator $\mathcal{O}_{t\Phi}$.
@@ VI Summary and Conclusions
In this study, we demonstrated that the GE representation of EW interactions, which we had a brief introduction first, has gauge symmetry imprinted in the structure of amplitudes, manifesting in the MWI (1). This approach to gauge symmetry has been rarely studied before, because it involves the vertices of both gauge bosons and Goldstone bosons simultaneously. We numerically studied this important property of EW interactions in the GE representation in several different aspects, including directly testing the MWI, modifying couplings within vertices, and modifying the overall couplings of vertices. Our main conclusion is that there are precise relations within and without the overall vertices, which guarantee the MWI. Any violation of these relations results in the violation of the MWI and gauge symmetry.
@@
By directly testing gauge symmetry on four-point amplitudes, such as $WW\rightarrow t\bar{t}$ and $W^{+}W^{-}\rightarrow W^{+}W^{-}$, with different helicity combinations, we observed that the full amplitudes always satisfy the MWI after summing over all tree-level diagrams. This result typically involves large cancellations between individual diagrams, similar to unitarity cancellation in the gauge representation.
@@
For testing gauge symmetry on three-point vertices with anomalous couplings, we studied the $VVh$, $ff^{\prime}V$, and $VVV$ vertices. We observed that there are precise relations governing the couplings of Goldstone and gauge components for these vertices. We then numerically tested if the MWI still holds when some of the couplings are modified so that those relations are violated. We found that any deviation from those relations by anomalous couplings would violate the MWI.
@@
For testing gauge symmetry on four-point amplitudes with anomalous couplings, we studied $^{+}W^{+}\rightarrow W^{+}W^{+}, WW$$W\rightarrow hh$, and $WW\rightarrow t\bar{t}$. We focused on how modifying the overall couplings of individual vertices affects gauge symmetry. Our results are similar to those for three-point vertices: gauge symmetry is manifested as precise relations between couplings, and modifying couplings to violate those relations also results in violating the MWI.
@@
After studying the gauge symmetry of the SM, we studied the effective operators in the SMEFT, specifically the $\mathcal{O}_{6}$ operator, which modifies the Higgs self-couplings, and the $\mathcal{O}_{t\Phi}$ operator, which modifies the top Yukawa coupling. We observed that the operators modify both the couplings mentioned and the related Goldstone couplings, giving relations between the couplings involving $C_{6}$ and $C_{t\Phi}$, respectively. Numerically testing the MWI indicates that gauge symmetry is preserved as long as all the related couplings are modified in accordance with those relations. On the other hand, if those relations are violated, gauge symmetry would be broken. Similar to the SM, our results for these SMEFT operators demonstrate that not only the gauge components but also the Goldstone components are crucial for maintaining gauge invariance of the theory. This is also the reason why, gauge cancellation in the unitary gauge, in which there is no Goldstone mode, appears to be ad hoc without an apparent physical mechanism.
@@
We believe that this paper contributes to a deeper understanding of EW interactions and massive gauge theory in general. In addition, it provides a convenient method of checking the self-consistency of EW interactions in the GE representation.
Thus far, the five-component formalism has been only systematically implemented in HELAS for the tree-level Feynman rules of the SM. Moreover, for the purpose of this study, we included the dim-6 operators $\mathcal{O}_{6}$ and $\mathcal{O}_{t\Phi}$ in HELAS. However, the majority of the SMEFT operators remain to be incorporated. This limits the utility of our method and calls for more progress in this aspect. In Ref. Hagiwara:2024xdh , a general method for automatically implementing the five-component framework has been proposed recently; which, we hope could help apply our method to more general theories.